A particular employee arrives at work sometime between 8:00 a.m. and 8:30 a.m. Based on past experience the company has determined that the employee is equally likely to arrive at any time between 8:00 a.m. and 8:30 a.m.

Find the probability that the employee will arrive before 8:07 a.m. Round your answer to four decimal places, if necessary.
Answer #1
here for uniform distribution parameter a =0 and b=30

probability that the employee will arrive before 8:07 a.m (t=7) :

P(X<7)=(7-0)/(30-0)=7/30 =0.2333